OpenStudy (anonymous):

What is the product written in scientific notation? (2.6 × 10–4)(2.2 × 108) A. 5.72 × 104 B. 5.72 × 10–32 C. 4.8 × 104 D. 4.8 × 10–2

OpenStudy (anonymous):

@AnswerMyQuestions @iGreen @I_Always_Smiling @Demonx341

OpenStudy (anonymous):

you think?

OpenStudy (anonymous):

Multiply 2.6 * 2.2 @thebro55

OpenStudy (anonymous):

@thebro55 Can you do that?

OpenStudy (anonymous):

\((2.6 \times 10^{–4})(2.2 \times 10^8) \rightarrow (2.6 \times 2.2) \times 10^{-4 + 8}\) Multiply 2.6 * 2.2, then add -4 + 8

OpenStudy (anonymous):

@thebro55 Are you there?